OpenStudy (anonymous):

Mark and Jaymee are 20 m apart. If they start running towards each other, Mark at 5.5 m/s and Jaymee at 4.5 m/s, what distance has mark covered by the time he meets jaymee?

Parth (parthkohli):

If Mark meets at x, then Jaymee would be at 20 - x. Time = Distance/Speed Time taken by mark to cover x km = x/5.5 ----By Jaymee = (20 - x)/4.5 The time that they'll meet at is equal. x/5.5 = 20-x/4.5 4.5x = 5.5(20 - x) 4.5x = 110 - 5.5x 10x = 110 x = 11 Mark has covered x meters = 11 meters Jaymee has covered 20 - x meters = 20 - 11 meters= 9 meters

OpenStudy (anonymous):

 left distant=5.5 m/s * t right distant= 4.5 m/s *t left +right =20 so 4.5t+5.5t=20 t=2 mark is the left one =2*5.5=11m

OpenStudy (anonymous):

Woahhhh. Hahaha. Genius XD Thanks! @IMMORTALS Hahaha, thanks for the effort :) You're a genius too! XD Thanks! :)